OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 12 and the last term is -3,359,232?

OpenStudy (campbell_st):

find the common ratio \[r ^{7} = -3359232/12\] then \[r = \sqrt[7]{(-339232/12)}\] then a = 12 and n = 8 and r was found above \[s _{n} = (a(r^n - 1)/(r -1))\]